• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 03:59:58 on June 8,2004

    Entry number 124938

    This entry is a followup to: 124932

    keyword=detector check(2) - note T0 change !!

    From a long cosmic run 1041. Summary: Detectors look ok,
    with two exceptions:

    1) L-arm S2 TDCs -- several channels are completely empty.
    2) R-arm S2 TDCs -- several channels are nearly empty.

    This is because the time offset changed and went out of TDC range.
    **NOTE** This may affect tracking !!
    Looking at VDC TDCs I see the shifts from previous experiment:
    L-arm T0 went from 1750 to now 1900. That is 150 chan = 75 nsec
    R-arm T0 went from 1800 to now 1850. That is 50 chan = 25 nsec.
    The L-arm shift seems a bit too large. Maybe something wrong ?
    Note that retiming signal (RT) is (and should be) unplugged.
    The ADC gate timings on both HRS were checked & ok today.
    The ADC gate is also wider than before: 300+ nsec.
    We can probably live without S2 timing info, but it might be better
    to move these into a TDC 1877 with bigger dynamic range (2 usec).

    In spite of the low efficiency due to the triggers, I can see
    data in S2, VDC, and leadglass. There was a small problem at
    first that D.evtypebits was zero. This was due to a timing
    shift which was slightly outside the cut window. I adjusted this
    in THaDynData. I have also updated the database file db_P.dat
    (note the "P"). This is for the ParityData class. The ADC and
    TDC locations for the HAPPEX detectors is set. See fig 1 for
    ADCs and TDCs for the HAPPEX e-cal detector.

    FIGURE 1